1.express each of these relationships as a rate in it's simplest form.
1.1. 12kg for R5.
1.2. 120 litres for 1000km.
1.3. R355 for two nights.
1.4. 7km in 28 min.
1.5. 145 students for five teachers.
1
Expert's answer
2020-10-06T16:28:55-0400

1.1 12 kg for Rs.5

Price for 1 kg = Rs. (5/12) = Rs. 0.4167 = Rs. 0.412

"\therefore" 1 kg for Rs.0.412


1.2 120 litres for 1000 km

Kms per litre = (1000/120) = 8.33 km

"\therefore" 1 litre for 8.33 km


1.3 Rs. 355 for two nights

Price per night = Rs. (355/2) = Rs. 177.5

"\therefore" Rs. 177.5 for one night


1.4 7 km in 28 min

Km in 1 mins = (7/28) km = 0.25 km

"\therefore" 0.25 km in 1 min


1.5 145 students for five teachers

No. Of Students for 1 teacher = 145/5 = 29

"\therefore" 29 students for 1 teacher